OpenStudy (chillhill):

f(x)= 1-log(base 5)(x-2), x>2 solve f(x)= -1 algebraically

OpenStudy (daniel.ohearn1):

Hey, so to do this I would use natural logs. Do you know to use ln?

OpenStudy (daniel.ohearn1):

Begin by setting it up as the equation -1=1-(logbase5ofx-2), x>2

OpenStudy (daniel.ohearn1):

logbase5of(x-2) = -2 implies logbase5 of x-2 = 2

OpenStudy (daniel.ohearn1):

Rewritten this means 5^y=2 where y=x-2.. From there take the natural log of each side. Alternatively you may take the log base y of both sides, before finally solving for x.